OpenStudy (phebe):

Which equation has a solution of ½? A. 13x – 12 = 14 B. 9x + 15 = 20 C. -6x – 18 = -21 D. -11x = 12x + 12

OpenStudy (anonymous):

C.

OpenStudy (phebe):

u sure

OpenStudy (phebe):

@pdxtwinz

OpenStudy (abhisar):

What do ub think @Phebe ?

OpenStudy (abhisar):

-6x – 18 = -21 \(\sf \Rightarrow-6x=-21+18\\ \Rightarrow6x=3\\ \Rightarrow x= \frac{3}{6}= \frac{1}{2}\)

OpenStudy (phebe):

thanks XD

OpenStudy (abhisar):

ur welcome :)

OpenStudy (phebe):

XD